OpenStudy (anonymous):

composition of functions: f(x)=3x-2, g(x)=x²+1, find (f°g)(-1) and (f°g)(3). can anyone help me understand how to go about it?

OpenStudy (anonymous):

fog means in f(x) put in g(x) in place of x

OpenStudy (anonymous):

yes first of all get rid of the circles. \[f\circ g(x)=f(g(x))\]

OpenStudy (anonymous):

so f(x) = 3x-2 and so f(g(x)) would be 3g(x) - 2

OpenStudy (anonymous):

now put in the definition of g(x)

OpenStudy (anonymous):

I'm so lost.

OpenStudy (anonymous):

then work from the inside out. start by replacing (x) by what it actually is in this case \[f(g(x)) =f(x^2+1)\] and now replace the x in f(x) by \[x^2+1\] to get \[3(x^2+1)-2\] \[3x^2+3-2\] \[3x^2+1\]

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

lets start with \[f\circ g(-1)\]

OpenStudy (anonymous):

what is \[g(-1)\]?

OpenStudy (anonymous):

is it clear how to find \[g(-1)\]?

OpenStudy (anonymous):

no its like I'm looking at a foreign language. I don't have any idea what I am supposed to be doing

OpenStudy (anonymous):

ok so lets do that first

OpenStudy (anonymous):

\[g(x)=x^2+1\] so if i want to find g of some number, i replace x by that number

OpenStudy (anonymous):

Yes functions I understand.

OpenStudy (anonymous):

\[g(-1)=(-1)^2+1=1+1=2\] \[g(2)=2^2+1=4+1=5\]

OpenStudy (anonymous):

ok so let me ask again, what is \[g(-1)\]?

OpenStudy (anonymous):

oh actually i wrote it for you. \[g(-1)=2\] yes?

OpenStudy (anonymous):

im a lost cause i guess

OpenStudy (anonymous):

so the only additional step to finding \[f\circ g(-1)\] is finding \[f(2)\] because that is what it means. first find g(-1) then find f of the result

OpenStudy (anonymous):

lets go back for a second. \[g(x)=x^2+1\] what is \[g(3)\]?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

of course this is not a lost cause. once you get it you will be like 'oh is that all?'

OpenStudy (anonymous):

right. now what is \[f(10)\]?

OpenStudy (anonymous):

i forgot what f was \[f(x)=3x-2\]

OpenStudy (anonymous):

28

OpenStudy (anonymous):

so now the question is what is \[f(10)\]

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

f(g(x) = f(x^2 +1) = 3(x^2+1) -2 = 3x^2 +1 fog(-1) = 3 (-1)^2 + 1 = 4 fog(3) = 3 *3^2 + 1 = 28

OpenStudy (anonymous):

so now we just have to understand what \[f \circ g(3)\] means. first you found that \[g(3)=10\] then you found that \[g(10)=28\] you are done.

OpenStudy (anonymous):

sorry i meant \[f(10)=28\]

OpenStudy (anonymous):

typo there. so the idea is a) first find g(3) b) then find f of the result. you have it

OpenStudy (anonymous):

but that's not the answer I should have. my text book tells me that it should be g(-1)=4 and f(3)=50

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

\[g(x)=x^2+1\] yes?

OpenStudy (anonymous):

no, first you calculate the fog formula and THEN the fog(3). etc

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so there is some problem with your text book not with you. unless you are looking at the wrong question

OpenStudy (anonymous):

it shows me step by step how it is solved. I'm not wrapping my mind around it though thats why I'm here

OpenStudy (anonymous):

because if \[g(x)=x^2+1\] then i guarantee you that \[g(-1)=(-1)^2+1=1+1=2\] you can bank on it

OpenStudy (anonymous):

first off we have to agree that \[g(-1)=2\] yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no matter what the text says.

OpenStudy (anonymous):

then it puts the 2 into the fog problem: 3(2)-2=4

OpenStudy (anonymous):

so now we want to compute \[f \circ g(-1)=f(g(-1))\]

OpenStudy (anonymous):

f(3) is where the problem is different

OpenStudy (anonymous):

not sure what you mean. \[f(3)=3\times 3 - 2=9-2=7\] but you are not asked for \[f(3)\]

OpenStudy (anonymous):

looks like you are asked for \[f(g(3))\] unless that is a typo and you want \[g(f(3))\]

OpenStudy (anonymous):

if the problem is written f(x)=3x-2, g(x)=x°+1 find fog (-1) and fog (3). why does it seem to be inputting them backwards? why does g(x) get he -1 and not the 3

OpenStudy (anonymous):

aaaaaaaaaaaaaaaahhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

no g(f(3)) doesn't give 50 i tried this

OpenStudy (anonymous):

they are giving you the answer to \[g(f(3))\]

OpenStudy (anonymous):

no it is not 50

OpenStudy (anonymous):

i feel so stupid why cant i understand this

OpenStudy (anonymous):

\[f(3)=7\] \[g(7)=7^2+1=50\] so \[g(f(3))+50\]

OpenStudy (anonymous):

you are not being stupid the text is wrong!

OpenStudy (anonymous):

either the text is wrong or the question are written wrong.

OpenStudy (anonymous):

lets make sure we have the problem correct with f and g correct as well. \[f(x)=3x-2\] \[g(x)=x^2+1\] for sure yes?

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

and not the other way around

OpenStudy (anonymous):

ok. and for sure you are asked for \[f\circ g(-1)\] not \[g \circ f(-1)\] yes? not the other way

OpenStudy (anonymous):

fog(-1) gof(3)

OpenStudy (anonymous):

ok \[f\circ g(-1)=f(g(-1)=f(2)=3\times 2-2=6-2=4\]

OpenStudy (anonymous):

ok now i see the problem. second question is \[g\circ f(3)\] not \[f\circ g(3)\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

they are different. for \[g\circ f(3)\] you need to compute \[f(3)\] first

OpenStudy (anonymous):

so first you compute \[f(3)\] and get \[g(3)=3\times 3-2=9-2=7\] yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

damn that was a typo. i mean \[f(3)=7\]

OpenStudy (anonymous):

ok and now you compute \[g(7)\] to get \[g(7)=7^2+1=49+1=50\] yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

whew. so we have \[g\circ f(3)=g(f(3))=g(7)=50\] and now the answers are correct and the same as in the text yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

idea is this \[g\circ f(x)\] means first do f, then do g whereas \[f\circ g(x)\] means first do g, then do f. they are different and almost always give different answers

OpenStudy (anonymous):

OMG that is the first thing to totally make sense to me

OpenStudy (anonymous):

its backwards (to me) but I can work with that

OpenStudy (anonymous):

that is because we read from left to right but we read functions the other way, from right to left

OpenStudy (anonymous):

you would thing \[f\circ g\] means first f then g because that is the way we read, but it does not

OpenStudy (anonymous):

the x is on the right we read \[f\circ g(x)\] means first \[g(x)\] then \[f(g(x))\] in other words first g, then f

OpenStudy (anonymous):

ahh. my text doesn't explain anything at all. I'm taking my class online so I cant raise my hand in class either

OpenStudy (anonymous):

thank you so much.I think I can try a few on my own

OpenStudy (anonymous):

would you mind checking my work on one?

OpenStudy (anonymous):

ok you can give me an example if you like i will be back in 5 minutes. i'll leave this open

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

just post it here

OpenStudy (anonymous):

f(x)=x°-2, g(x)=x+4 find fog (2) and gof (-4) g(x)=x+4 g(s)= 2+4=6 f(-4)=(-4)°-2 =16-2=14 fog(6)=f(g(6)) =6°-2=36-2=34 gof (14)=g(f(14)) 14+4=18

OpenStudy (anonymous):

ok let me check \[f(x)=x^2-2\] \[g(x)=x+4\] \[f\circ g(2)=f(g(2))=f(6)=6^2-2=36-2=34\]

OpenStudy (anonymous):

looks good to me!

OpenStudy (anonymous):

\[g\circ f(-4)=g(f(-4))=g(14)=14+4=18\]

OpenStudy (anonymous):

that one looks good too!

OpenStudy (anonymous):

wow i think I understand it

OpenStudy (anonymous):

you are a lifesaver

OpenStudy (anonymous):

whew. not to say "i told you so" but \[\color{red}{\text{i told you so!}}\]

OpenStudy (anonymous):

not that hard once you know what you are doing. now before you go

OpenStudy (anonymous):

lets make sure you can also compute \[f\circ g(x)\] in the above example

OpenStudy (anonymous):

lol. thank you thank you thank you!!!!

OpenStudy (anonymous):

we do it the same way, but with "x" instead of numbers.

OpenStudy (anonymous):

\[f\circ g(x)=f(g(x))=f(x+4)=(x+4)^2-2\]

OpenStudy (anonymous):

and you see if you replace x by 2 in the above you will get \[f \circ g(2))=f(g(2))=(2+4)^2-2=36-2=34\] just like before